OpenStudy (anonymous):

Help me out people. Medal awarded and I'll love you forever. The present value that will give a future value S in 3 years with interest compounded annually varies inversely as the cube of 1+r, where r is the annual interest rate. If present value of $8213.19 gives a future value of 10000, what would be the present value of 16600?

OpenStudy (anonymous):

(8213.19)[(1 + r)^3] = 10000 (1 + r)^3 = 10000/8213.19 (PV)[(1 + r)^3] = 16600 -> PV = 16600/[(1 + r)^3] PV = 16600 / (10000/8213.19)

OpenStudy (anonymous):

All good now, @jmallis ?

OpenStudy (anonymous):

Yes thank-you so much

OpenStudy (anonymous):

uw! Good luck in all of your studies and thx for the recognition! @jmallis

OpenStudy (anonymous):

Thank-you again